# Supervised Graph Contrastive Learning for Gene Regulatory Networks

### v4

@@ Abstract
Graph Contrastive Learning (GCL) is a powerful self-supervised learning framework that performs data augmentation through graph perturbations, with growing applications in the analysis of biological networks such as Gene Regulatory Networks (GRNs). The artificial perturbations commonly used in GCL, such as node dropping, induce structural changes that can diverge from biological reality. This concern has contributed to a broader trend in graph representation learning toward augmentation-free methods, which view such structural changes as problematic and to be avoided. However, this trend overlooks the fundamental insight that structural changes from biologically meaningful perturbations are not a problem to be avoided but a rich source of information, thereby ignoring the valuable opportunity to leverage data from real biological experiments. Motivated by this insight, we propose SupGCL (Supervised Graph Contrastive Learning), a new GCL method for GRNs that directly incorporates biological perturbations from gene knockdown experiments as supervision. SupGCL is a probabilistic formulation that continuously generalizes conventional GCL, linking artificial augmentations with real perturbations measured in knockdown experiments and using the latter as explicit supervisory signals. To assess effectiveness, we train GRN representations with SupGCL and evaluate their performance on downstream tasks. The evaluation includes both node-level tasks, such as gene function classification, and graph-level tasks on patient-specific GRNs, such as patient survival hazard prediction. Across 13 tasks built from GRN datasets derived from patients with three cancer types, SupGCL consistently outperforms state‑of‑the‑art baselines.
@@ 1 Introduction
However, this trend overlooks a fundamental insight: that the representational shifts caused by graph augmentation are not an obstacle to overcome, but rather a rich source of information to be exploited. Consequently, it ignores the valuable opportunity to leverage data obtained from actual biological experiments, such as knockdown experiments.
To address these issues, we propose a novel supervised GCL method (SupGCL) that leverages gene knockdown perturbations within GRNs. Our method uses experimental data from actual gene knockdowns as supervision, enabling biologically faithful representation learning. In gene knockdown experiments, the suppression of specific genes causes biological perturbations, leading to the observation of a new, altered GRN. By using these perturbations as supervision signals for GCL, we can perform data augmentation that retains biological characteristics. Moreover, since our method naturally extends traditional GCL models in the direction of supervised augmentation within a probabilistic framework, conventional GCL approaches emerge as special cases of our proposed model.
To evaluate the effectiveness of the proposed SupGCL method, we apply it to GRN datasets from cancer patients across three cancer types and conduct multiple downstream tasks. For gene-level downstream tasks, we perform classification into biological process, cellular component, and cancer-related gene categories. For patient-level tasks, we conduct hazard prediction and disease subtype classification. The performance of our method is compared against existing graph representation learning techniques, including conventional GCL methods.
The main contributions of this study are as follows:
  • Proposal of a novel GRN representation learning method utilizing gene knockdown experiments: We develop a new GCL method tailored for GRNs that incorporates gene knockdown data as supervision to enhance biological plausibility.
  • Theoretical extension of GCL: We formulate supervised GCL, incorporating augmentation selection into a unified probabilistic modeling framework, and theoretically demonstrate that existing GCL methods are special cases of our proposed approach.
  • Empirical validation of the proposed method: We apply the method to 13 downstream tasks on GRNs derived from real cancer patients and consistently outperform conventional approaches across all tasks.
@@ 2 Related Works
2.1 Augmentation-Free Graph Contrastive Learning (GCL)
GCL is a powerful framework for learning representations by maximizing the similarity between multiple ”views” of a graph generated through data augmentation. While early node-level methods like GRACE achieved high performance (Zhu et al., 2020), graph-level approaches like GraphCL struggled with the loss of key node features (You et al., 2021; Sun et al., 2025).
A common challenge for these methods was their heavy reliance on the choice of data augmentations. This led to a trend in augmentation-free GCL, pioneered by BGRL (Thakoor et al., 2021), which used a bootstrapping mechanism to circumvent heuristic augmentations that could damage the graph’s structure, influencing subsequent research such as simGRACE (Xia et al., 2022) and AFGRL (Lee et al., 2022). SGRL (He et al., 2024) further advanced this trend by integrating concepts like feature uniformity from early GCL to prevent the representation collapse problem often faced by methods like BGRL, achieving stable and high-performance self-supervised learning. However, these methods view perturbations to the graph as a problem to be avoided, thereby overlooking the opportunity to utilize real-world supervisory signals that arise from structural changes.
2.2 Representation Learning for Gene Regulatory Networks(GRNs)
@@
Applying Graph Neural Networks (GNNs) to known GRNs is a growing area of research for downstream tasks such as gene function classification and patient survival prediction(Zohari & Chehreghani, 2025). The prevailing approach in this field has been supervised learning without pre-training (Liu et al., 2022; Yu et al., 2025). While link prediction addresses the important task of inferring the unknown structure of GRNs (Huynh-Thu et al., 2010; Yu et al., 2025), our work focuses on learning representations from known GRN structures for downstream applications.
@@
Recently, self-supervised methods have been introduced to this domain. Initial efforts focused on applying general augmentation-free methods, such as the Graph Autoencoder (GAE), which learns representations by reconstructing the graph’s structure (Jung et al., 2024). More recently, the focus has begun to shift towards specialized GCL frameworks tailored for GRNs, although such approaches are still uncommon. A notable example is MuSe-GNN (Liu et al., 2023), which leverages different data modalities (e.g., scRNA-seq and scATAC-seq) as natural ”views,” avoiding the need for artificial augmentations by using biologically diverse information. In contrast to these methods, our work, SupGCL, uniquely incorporates real-world perturbations from within a single modality as supervisory signals for GCL.
@@ 3.1 Background of Graph Contrastive Learning
$p(j|i)\triangleq\delta_{ij},\quad q^{a,b}_{\phi}(j|i)\triangleq\frac{\exp(\mathrm{sim}({\bm{z}}_{i}^{a},{\bm{z}}_{j}^{b})/\tau_{\rm n})}{\sum_{k\in\mathcal{V}}\exp(\mathrm{sim}({\bm{z}}_{i}^{a},{\bm{z}}_{k}^{b})/\tau_{\rm n})}.$ (2)
@@
While GCL achieves node-level representation learning via the procedure described above, in many cases the augmentation operations themselves rely on artificial perturbations such as randomly adding and/or deleting nodes and/or edges. In this study, we introduce gene knockdown—a biological perturbation—as supervision for these augmentation operations.
@@ 3.2 Notation and Problem Definition
${\bm{Z}}^{a}\triangleq f_{\phi}(\mathcal{G}_{a})\in\mathbb{R}^{|\mathcal{V}|\times d},\quad{\bm{Y}}^{a}\triangleq f_{\phi}(\mathcal{H}_{a})\in\mathbb{R}^{|\mathcal{V}|\times d},$ (5)
@@ 4 Method
For the set of embedded representations $\{({\bm{Z}}^{a},{\bm{Y}}^{a})\}_{a\in\mathcal{K}}$, we consider the pair of augmentation operations $(a,b)$ and the pair of nodes $(i,j)$ according to the contractive learning scheme. First, for the pair of augmentation operations $(a,b)$, we clarify the supervised learning problem for augmentation operations using KL divergence and then propose SupGCL using a distribution over pairs of combinations of nodes and extension operations. A sketch of the proposed method is shown in Figure 1.
@@
$p_{\phi}(b|a)\triangleq\frac{\exp\left(\mathrm{sim}_{F}({\bm{Y}}^{a},{\bm{Y}}^{b})/\tau_{\rm a}\right)}{\sum_{c\in\mathcal{K}}\exp\left(\mathrm{sim}_{F}({\bm{Y}}^{a},{\bm{Y}}^{c})/\tau_{\rm a}\right)},\quad q_{\phi}(b|a)\triangleq\frac{\exp\left(\mathrm{sim}_{F}({\bm{Z}}^{a},{\bm{Z}}^{b})/\tau_{\rm a}\right)}{\sum_{c\in\mathcal{K}}\exp\left(\mathrm{sim}_{F}({\bm{Z}}^{a},{\bm{Z}}^{c})/\tau_{\rm a}\right)},$ (6)
@@
Minimizing this loss reduces the discrepancy in embedding distributions between the artificially augmented graphs and the biologically grounded knockdown graphs. However, if both $p_{\phi}$ and $q_{\phi}$ are optimized simultaneously, the model may converge to a trivial solution. For instance, if the GNN outputs constant embeddings, both distributions become uniform and $\mathrm{Loss}_{\rm Aug}=0$. Thus, minimizing $\mathrm{Loss}_{\rm Aug}$ alone is insufficient for learning meaningful graph representations.
To address this issue, here we first introduce a reference model $p_{\phi}(j,b|i,a)$ and a learned model $q_{\phi}(j,b|i,a)$ that use conditional probabilities for each pair of node and augmentation $(i,a),(j,b)\in\mathcal{V}\times\mathcal{K}$. By substituting these into the contrastive learning formulation in (1), we derive the loss function of Supervised Graph Contrastive Learning:
$\mathrm{Loss}_{\rm SupGCL}\triangleq\frac{1}{|\mathcal{V}||\mathcal{K}|}\sum_{i\in\mathcal{V},a\in\mathcal{K}}D_{\rm KL}(p_{\phi}(j,b|i,a)|q_{\phi}(j,b|i,a)).$ (8)
@@
$\mathrm{Loss}_{\rm SupGCL}=\mathbb{E}_{a,b\sim p_{\phi}(b|a)\mathrm{U}_{\mathcal{K}}(a)}\big[{\rm Loss}_{\rm node}^{a,b}\big]+\mathrm{Loss}_{\rm Aug}.$ (9)
Proof:.
This follows directly from the standard decomposition of KL divergence: $D_{\rm KL}(p(x,y)|q(x,y))=\mathbb{E}_{x\sim p(x)}[D_{\rm KL}(p(y|x)|q(y|x))]+D_{\rm KL}(p(x)|q(x))$. See Appendix A for details. ∎
The first term in Theorem 1 corresponds to the expectation of the node-level GCL loss $\mathrm{Loss}_{\rm node}^{a,b}$ (as defined in (3)) with respect to the supervised augmentation distribution $p_{\phi}(b|a)$. This allows node-level contrastive learning to reflect biological similarity between knockdown operations. Importantly, since the theorem is independent of the specific choice of the node-level model $(p,q_{\phi}^{a,b})$, any contrastive loss described by KL divergence can be used in practice. Meanwhile, the second term reduces the distributional difference between the artificially generated augmentation-based GRN and the teacher GRN. Together, these two components ensure both expressive node representations and biologically meaningful augmentations.
Moreover, the performance of node-level representation learning and the biological validity following the teacher data for augmentation operations can be controlled by the temperature parameters $\tau_{\rm n}$ and $\tau_{\rm a}$ of each probability model. In particular, when the temperature parameter $\tau_{\rm a}$ involved in the augmentation operation is sufficiently large, the augmentation operation becomes independent of the teacher GRNs $\{{\bm{Y}}_{a}\}_{a\in\mathcal{K}}$, and coincides with the conventional node-level GCL loss function.
Corollary 1.
@@
Proof:.
As $\tau_{\rm a}\rightarrow\infty$, we have $p_{\phi}(b|a)\rightarrow\mathrm{U}_{\mathcal{K}}(b)$ and $q_{\phi}(b|a)\rightarrow\mathrm{U}_{\mathcal{K}}(b)$. Therefore, the expectation term becomes: $\lim_{\tau_{\rm a}\rightarrow\infty}\mathbb{E}_{a,b\sim p_{\phi}(b|a)\mathrm{U}_{\mathcal{K}}(a)}\big[{\rm Loss}_{\rm node}^{a,b}\big]=\mathbb{E}_{a,b\sim{\rm U}_{\mathcal{K}}}[\mathrm{Loss}_{\rm node}^{a,b}]$, $\lim_{\tau_{\rm a}\rightarrow\infty}D_{\mathrm{KL}}(p_{\phi}(b|a)|q_{\phi}(b|a))=0$ thus proving the corollary. ∎
In this study, we train the GNN using standard gradient-based optimization applied to the loss function defined in Theorem 1. The corresponding pseudocode is provided in Appendix B.

### v5

@@ Abstract
Graph Contrastive Learning (GCL) is a powerful self-supervised learning framework that performs data augmentation through graph perturbations, with growing applications in the analysis of biological networks such as Gene Regulatory Networks (GRNs). The artificial perturbations commonly used in GCL, such as node dropping, induce structural changes that can diverge from biological reality. This concern has contributed to a broader trend in graph representation learning toward augmentation-free methods, which view such structural changes as problematic and should be avoided. However, this trend overlooks the fundamental insight that structural changes from biologically meaningful perturbations are not a problem to be avoided, but rather a rich source of information, thereby ignoring the valuable opportunity to leverage data from real biological experiments. Motivated by this insight, we propose SupGCL (Supervised Graph Contrastive Learning), a new GCL method for GRNs that directly incorporates biological perturbations from gene knockdown experiments as supervision. SupGCL is a probabilistic formulation that continuously generalizes conventional GCL, linking artificial augmentations with real perturbations measured in knockdown experiments, and using the latter as explicit supervision. On patient-derived GRNs from three cancer types, we train GRN representations with SupGCL and evaluate it in two regimes: (i) embedding space analysis, where it yields clearer disease-subtype structure and improves clustering, and (ii) task-specific fine-tuning, where it consistently outperforms strong graph representation learning baselines on 13 downstream tasks spanning gene-level functional annotation and patient-level prediction.
Machine Learning, ICML
[…]
@@ 1 Introduction
However, this trend overlooks a fundamental insight: that the representational shifts caused by graph augmentation are not an obstacle to overcome but rather a rich source of information to be exploited. Consequently, it ignores the valuable opportunity to leverage data obtained from actual biological experiments, such as knockdown experiments, even though they have become readily accessible enabled by advances in high-throughput profiling technologies (Subramanian et al., 2017a).
To address these issues, we propose a novel supervised GCL method (SupGCL) that leverages gene knockdown perturbations within GRNs. Our method uses experimental data from actual gene knockdowns as supervision, enabling biologically faithful representation learning. In gene knockdown experiments, the suppression of specific genes causes biological perturbations, leading to the observation of a new altered GRN. By using these perturbations as supervision signals for GCL, we can perform data augmentation that retains biological characteristics. Moreover, since our method naturally extends traditional GCL models in the direction of supervised augmentation within a probabilistic framework, node-level GCL approaches emerge as special cases of our proposed model.
To evaluate the effectiveness of the proposed method, we benchmark SupGCL on patient-derived GRNs from three cancer types in both embedding space analysis and fine-tuning. Without any task-specific training, SupGCL produces markedly better disease-subtype structure in the embedding space, improving clustering quality. With fine-tuning, it consistently surpasses strong graph representation learning baselines—including previous GCL models—across 13 downstream tasks spanning gene-level functional annotation (Gene Ontology categories and cancer-gene classes) and patient-level tasks (survival risk and subtype classification). Our main contributions are as follows:
  • Redefining Graph Augmentation by Introducing Biological Perturbations: We develop a new GCL method for GRNs that incorporates gene knockdown data as the supervision of graph perturbations to enhance biological plausibility.
  • Theoretical Extension of Node-level GCL: We formulate supervised GCL, incorporating augmentation selection into a unified probabilistic modeling framework, and theoretically demonstrate that existing node-level GCL methods correspond to singular solutions under this formulation.
  • Empirical Validation: We quantitatively demonstrate that SupGCL achieves higher clustering metrics (NMI/ARI) for disease subtypes without task-specific training. Furthermore, it consistently outperforms state-of-the-art baselines across 13 downstream tasks for three cancer types through fine-tuning.
@@ 2 Related Works
Augmentation-Free GCLs :
GCL learns representations by maximizing agreement between multiple augmented views of a graph. While early node-level methods such as GRACE achieved strong performance (Zhu et al., 2020), graph-level methods like GraphCL often suffered from losing key node features under augmentation (You et al., 2021; Sun et al., 2025). These methods remained a common limitation that is the heavy dependence on augmentation choices. This limitation motivated augmentation-free GCL, initiated by BGRL (Thakoor et al., 2021), which replaces heuristic (and potentially structure-damaging) augmentations with a bootstrapping mechanism, inspiring follow-up work such as simGRACE (Xia et al., 2022a) and AFGRL (Lee et al., 2022). SGRL (He et al., 2024) further incorporates principles such as feature uniformity to mitigate representation collapse observed in methods like BGRL, yielding stable and high-performing self-supervised learning. However, these approaches treat graph perturbations primarily as nuisances to avoid, and thus miss opportunities to exploit real-world supervisory data by structural changes.
Augmentation-Adjustment GCLs :
To reduce reliance on manually designed augmentations, another line of work learns augmentations by directly optimizing graph perturbations. In this setting, AD-GCL (Suresh et al., 2021) and ArieL (Feng et al., 2024) generate robust views via adversarial training, while AutoGCL (Yin et al., 2022) learns diverse augmentation policies that preserve the labels of contrastive pairs. By endogenizing augmentation rather than treating it as a fixed noise source, these frameworks support robust learning across diverse graph structures. However, since their perturbations are driven solely by optimization objectives, they do not effectively expand data with variations that reflect real-world phenomena.
Statistical Formulations of Contrastive Learning :
The theoretical formulation of contrastive learning has been actively evolving. The I-Con framework (Alshammari et al., 2025) presents a unified statistical formulation of representation learning by adopting a variational-inference perspective to organize contrastive learning methods as the minimization of KL divergences between an ideal reference distribution and a model-induced one. Within this framework, SupCon (Khosla et al., 2020), proposed as a supervised contrastive learning model, constructs the reference distribution using label data to concentrate features of the same class. However, such methods fundamentally rely on one-to-one supervised categorical labels for each sample and remain restricted to probabilistic modeling at the sample level.
Representation Learning for GRNs :
@@
Applying Graph Neural Networks to known GRNs is a growing area of research for downstream tasks such as gene function classification and patient survival prediction(Zohari & Chehreghani, 2025). The prevailing approach in this field has been supervised learning without pre-training (Liu et al., 2022; Yu et al., 2025). While link prediction addresses the important task of inferring the unknown structure of GRNs (Huynh-Thu et al., 2010; Yu et al., 2025), our work focuses on learning representations from known GRN structures for downstream applications.
@@
Recently, self-supervised methods have been introduced to this domain. Initial efforts focused on applying general augmentation-free methods, such as the Graph Autoencoder, which learns representations by reconstructing the graph’s structure (Jung et al., 2024). More recently, the focus has begun to shift towards specialized GCL frameworks tailored for GRNs, although such approaches are still uncommon. A notable example is MuSe-GNN (Liu et al., 2023), which leverages different data modalities (e.g., scRNA-seq and scATAC-seq) as natural ”views,” avoiding the need for artificial augmentations by using biologically diverse information. In contrast to these methods, our work, SupGCL, uniquely incorporates real-world perturbations from within a single modality as supervisory signals for GCL.
@@ 3.1 Background of Graph Contrastive Learning
$p(j|i)\triangleq\delta_{ij},\quad q^{a,b}_{\phi}(j|i)\triangleq\frac{\exp(\mathrm{sim}({\bm{z}}_{i}^{a},{\bm{z}}_{j}^{b})/\tau_{\rm n})}{\sum_{k\in\mathcal{V}}\exp(\mathrm{sim}({\bm{z}}_{i}^{a},{\bm{z}}_{k}^{b})/\tau_{\rm n})}.$
@@
While GCL achieves node-level representation learning via the procedure described above, in many cases the augmentation operations themselves rely on artificial perturbations such as randomly adding and/or deleting nodes and/or edges. In this study, we introduce gene knockdown – a biological perturbation – as a supervision for these augmentation operations.
@@ 3.2 Notation and Problem Definition
${\bm{Z}}^{a}\triangleq f_{\phi}(\mathcal{G}_{a})\in\mathbb{R}^{|\mathcal{V}|\times d},\quad{\bm{Y}}^{a}\triangleq f_{\phi}(\mathcal{H}_{a})\in\mathbb{R}^{|\mathcal{V}|\times d},$
@@ 4 Method
For the set of embedded representations $\{({\bm{Z}}^{a},{\bm{Y}}^{a})\}_{a\in\mathcal{K}}$, we consider the pair of augmentation operations $(a,b)$ and the pair of nodes $(i,j)$ according to the contrastive learning scheme. First, for the pair of augmentation operations $(a,b)$, we clarify the supervised learning problem for augmentation operations using KL divergence and then propose SupGCL using a distribution over pairs of combinations of nodes and extension operations. A sketch of the proposed method is shown in Figure 1.
@@
$p_{\phi}(b|a)$ $\triangleq\frac{\exp\left(\mathrm{sim}_{F}({\bm{Y}}^{a},{\bm{Y}}^{b})/\tau_{\rm a}\right)}{\sum_{c\in\mathcal{K}}\exp\left(\mathrm{sim}_{F}({\bm{Y}}^{a},{\bm{Y}}^{c})/\tau_{\rm a}\right)},$
$q_{\phi}(b|a)$ $\triangleq\frac{\exp\left(\mathrm{sim}_{F}({\bm{Z}}^{a},{\bm{Z}}^{b})/\tau_{\rm a}\right)}{\sum_{c\in\mathcal{K}}\exp\left(\mathrm{sim}_{F}({\bm{Z}}^{a},{\bm{Z}}^{c})/\tau_{\rm a}\right)},$
@@
Minimizing this loss reduces the discrepancy in embedding distributions between artificially augmented graphs and biologically grounded knockdown graphs. However, if both $p_{\phi}$ and $q_{\phi}$ are optimized simultaneously, the model may converge to a trivial solution. For example, if the GNN outputs constant embeddings, both distributions become uniform, and $\mathrm{Loss}_{\rm Aug}=0$. Thus, minimizing $\mathrm{Loss}_{\rm Aug}$ alone is insufficient for learning meaningful graph representations.
To address this issue, here we first introduce a reference model $p_{\phi}(j,b|i,a)$ and a learned model $q_{\phi}(j,b|i,a)$ that use conditional probabilities for each pair of nodes and augmentation $(i,a),(j,b)\in\mathcal{V}\times\mathcal{K}$. By substituting these into the contrastive learning formulation in (1), we derive the loss function of Supervised Graph Contrastive Learning:
$\mathrm{Loss}_{\rm SupGCL}$
$\triangleq\frac{1}{|\mathcal{V}||\mathcal{K}|}\sum_{i\in\mathcal{V},a\in\mathcal{K}}D_{\rm KL}(p_{\phi}(j,b|i,a)|q_{\phi}(j,b|i,a)).$ (5)
@@
$\mathrm{Loss}_{\rm SupGCL}$
$=\mathbb{E}_{a,b\sim p_{\phi}(b|a)\mathrm{U}_{\mathcal{K}}(a)}\big[{\rm Loss}_{\rm Node}^{a,b}\big]+\mathrm{Loss}_{\rm Aug}.$ (6)
Proof.
This follows directly from the standard decomposition of KL divergence: $D_{\rm KL}(p(x,y)|q(x,y))=\mathbb{E}_{x\sim p(x)}[D_{\rm KL}(p(y|x)|q(y|x))]+D_{\rm KL}(p(x)|q(x))$. See Appendix A for more details. ∎
The first term in Theorem 4.1 corresponds to the expectation of the node-level GCL loss $\mathrm{Loss}_{\rm Node}^{a,b}$ (as defined in (2)) with respect to the supervised augmentation distribution $p_{\phi}(b|a)$. This allows node-level contrastive learning to reflect biological similarity between knockdown operations. Importantly, since the theorem is independent of the specific choice of the node-level model $(p,q_{\phi}^{a,b})$, any contrastive loss described by KL divergence can be used in practice. Meanwhile, the second term reduces the distributional difference between the artificially generated augmentation-based GRN and the teacher GRN. Together, these two components ensure both expressive node representations and biologically meaningful augmentations.
Moreover, the performance of the node-level representation learning and the biological validity following teacher data for the augmentation operations can be controlled by the temperature parameters $\tau_{\rm n}$ and $\tau_{\rm a}$ of each probability model. In particular, when the temperature parameter $\tau_{\rm a}$ involved in the augmentation operation is sufficiently large, the augmentation operation becomes independent of the teacher GRNs $\{{\bm{Y}}_{a}\}_{a\in\mathcal{K}}$, and coincides with the conventional node-level GCL loss function.
Corollary 4.2.
@@
Proof.
As $\tau_{\rm a}\rightarrow\infty$, we have $p_{\phi}(b|a)\rightarrow\mathrm{U}_{\mathcal{K}}(b)$ and $q_{\phi}(b|a)\rightarrow\mathrm{U}_{\mathcal{K}}(b)$. Therefore, the expectation term becomes $\lim_{\tau_{\rm a}\rightarrow\infty}\mathbb{E}_{a,b\sim p_{\phi}(b|a)\mathrm{U}_{\mathcal{K}}(a)}\big[{\rm Loss}_{\rm Node}^{a,b}\big]=\mathbb{E}_{a,b\sim{\rm U}_{\mathcal{K}}}[\mathrm{Loss}_{\rm Node}^{a,b}]$, $\lim_{\tau_{\rm a}\rightarrow\infty}D_{\mathrm{KL}}(p_{\phi}(b|a)|q_{\phi}(b|a))=0$ , thus proving the corollary. ∎
In this study, we train the GNN using standard gradient-based optimization applied to the loss function defined in Theorem 4.1. The corresponding pseudocode is provided in Appendix B.
Remark 4.3.
When supervision is available for graph augmentations, one may consider learning the augmentation function itself as a possible strategy. However, such an approach results in an ill-posed formulation within the contrastive learning framework admitting trivial solutions. An analysis of the validity of SupGCL in these augmentation settings, together with a quantitative comparison across different augmentation schemes, is provided in the Appendix. L.